OpenStudy (anonymous):

Given: cos(theta) = -4/5 , sin x = -12/13, theta is in the third quadrant, and x is in the fourth quadrant; evaluate tan(1/2)(theta) -3 3 1/3

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

So I know that tan = 3/4 (from the last problem) but do I just multiply that by 1/2? Because that would make it 3/8 but that's not a choice... What do I do?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

is this the question: \(\large tan(\frac{\theta}{2})= \) ????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok.. do you know the half-angle formula for tangent?

OpenStudy (anonymous):

nvmd... i'll put it up..

OpenStudy (anonymous):

\(\huge tan(\frac{\theta}{2})=\pm\sqrt{\frac{1-cos\theta}{1+cos\theta}} \)

OpenStudy (anonymous):

now you already have the value for cos(theta). what was it???

OpenStudy (anonymous):

Oh right right right!! I forgot the formulas part.. & cos(theta)=3/4

OpenStudy (anonymous):

the problem says cos(theta) = -4/5

OpenStudy (anonymous):

Whoops, that was tan.. Embarassing.

OpenStudy (anonymous):

sawright... :)

OpenStudy (anonymous):

So would that give you 9?

OpenStudy (anonymous):

so.... \(\large tan(\frac{\theta}{2})=\pm\sqrt{\frac{1-(-4/5)}{1+(-4/5)}} \) simplify this.... then we'll determine which one it is... (positive or negative)

OpenStudy (anonymous):

the top would be 9/5 and the bottom 1/5 then you would get 9 & you have to sqrt it. So it would be +/- 3